calculate the pH when 50.0 mL of 0.4 M NH3 is mixed with 5.00 mL of 0.8 M HCl. ( The ka of ammonia is 1.77 x10^-5.)


KaKb=Kw
(Ka)(Kb)=1.00 x 10^-14
Answer #1

Given:

M(HCl) = 0.8 M

V(HCl) = 5 mL

M(NH3) = 0.4 M

V(NH3) = 50 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.8 M * 5 mL = 4 mmol

mol(NH3) = M(NH3) * V(NH3)

mol(NH3) = 0.4 M * 50 mL = 20 mmol

We have:

mol(HCl) = 4 mmol

mol(NH3) = 20 mmol

4 mmol of both will react

excess NH3 remaining = 16 mmol

Volume of Solution = 5 + 50 = 55 mL

[NH3] = 16 mmol/55 mL = 0.2909 M

[NH4+] = 4 mmol/55 mL = 0.0727 M

They form basic buffer

base is NH3

conjugate acid is NH4+

Kb = 1.77*10^-5

pKb = - log (Kb)

= - log(1.77*10^-5)

= 4.752

use:

pOH = pKb + log {[conjugate acid]/[base]}

= 4.752+ log {7.273*10^-2/0.2909}

= 4.15

use:

PH = 14 - pOH

= 14 - 4.15

= 9.85

Answer: 9.85